Donaldson Mfg. is considering a capital budgeting project with a cost of capital of 10% and the following expected cash flow pattern:

Time 0 1 2 3 4 5
Cash flow -100 25 50 50 25 10

What is the profitabiity index of this project?

0.75

1.25

1.22

1.35
Answer #1

Present value = 25 / (1 + 0.1)1 + 50 / (1 + 0.1)2 + 50 / (1 + 0.1)3 + 25 / (1 + 0.1)4 + 10 / (1 + 0.1)5

Present value = 22.7273 + 41.3223 + 37.5657 + 17.0753 + 6.209213

Present value = 124.8998

Profitabiity index = Present value / initial cost

Profitabiity index = 124.8998 / 100

Profitabiity index = 1.25